bill533:

Is this statement true or false? The point (5, -1) is a solution of the equation: y = 2x − 11 A. true B. false

dude:

To check, just substitute the point into the equation, if theyre equal then its true, otherwise its false

jhonyy9:

bill do you understand what dude wrote above ?

jhonyy9:

this given point has coordinates x=5 and y = -1 so just substitute them inside this given equation and check that is true or not

jhonyy9:

-1 = 2*5 -11 is this true or false ?

bill533:

True

jhonyy9:

yes bc. -1 = 10-11 so -1 = -1

bill533:

Put "5" in the function for x......see if this returns "-1" If so....it is a solution

jhonyy9:

yes sure

bill533:

thanks

jhonyy9:

np with my pleasure anytime